choose the set that is the possible side lengths of a right triangle
a)4,9,13
b)squarer 2, squarer 2, 2
c)1,1,2
d)8,15,25

Answer by richwmiller(17219)

not a 4+9=13
not c 1+1=2
not d 8+15=23
Guess which